the sum of 1/3 and three times a number is equal to 4/5 subtracted from four times the number find the number. is it 7/15?
Answer by dabanfield(803)

the sum of 1/3 and three times a number is equal to 4/5 subtracted from four times the number find the number. is it 7/15?
Let x be the number.
Then we have:
1/3 + 3*x = 4*x - 4/5
4*x - 3*x = 4/5 + 1/3
x = 12/15 + 5/15 = 17/15